OpenStudy (anonymous):

6.2(-3x)+(4.1)(-3x) what do i do ? (:

OpenStudy (anonymous):

multiply and simplify

OpenStudy (anonymous):

multiply what ?

OpenStudy (anonymous):

so you muliply 6.2 times -3x you get -18.6x then you multiply the other which is 4.1 times -3x and you get -12.3x then add -18.6x and -12.3x and you get -30.9x

OpenStudy (cwrw238):

6.2 * -3x = -18.6x do second part in same way

OpenStudy (anonymous):

ohh okay thanks guys (:

OpenStudy (cwrw238):

and add them

OpenStudy (anonymous):

np(:

OpenStudy (anonymous):

You could use the distributive property; it might make the problem simpler\[6.2(-3x)+(4.1)(-3x)=-3x(6.2+4.1)=-3x(10.3)=-30.9x\]

OpenStudy (unklerhaukus):

factorize \[(6.2)(-3x)+(4.1)(-3x)=-3x(6.2+4.1)\]